OpenStudy (anonymous):

Determine the number of solutions you will have for the system 4y + x = 16 and y = 4 - x. Select one: a. No solutions. b. One solution. c. Impossible to determine. d. Indefinite solutions.

OpenStudy (anonymous):

\[4y+x = 16\]\[y = 4-x\] Since you have a system you can substitute your y into your first equation. \[4(4-x)-x = 16\] And from there you can solve for the value of x. Depending on the outcome you can answer your question.